The mean return for a random sample of 33 mutual funds is 14.93 percent with a standard deviation of 9.57. Test the null hypothesis: H0: mu = 15% at α = 0.05
1
Expert's answer
2011-06-14T09:59:47-0400
t-value=(x-15)/s*sqrt(n)=(14.93-15)/9.57*sqrt(33)= -0.042

p-value=P(T < t-value, with df=33-1=32 ) = 0.48 > 0.05
So p-value is more than alpha and we fail to reject the our null
hypotheis about m=15%